Find the slope of each line and the coordinates of one point on the line.
a) 2(y-2)=3(x+5)
b) y=-5+2(x+1)

Answer by stanbon(75887)

Find the slope of each line and the coordinates of one point on the line.
a) 2(y-2)=3(x+5)
2y-4 = 3x+15
2y = 3x+19
y = (3/2)x + (19/2)
slope = 3/2
y-int: (0,19/2)
--------------------
b) y=-5+2(x+1)
y = 2x+2-3
y = 2x-3
---------
slope = 2
y-int: (0,-3)
====================
Cheers,
Stan H.
===========